3. How many excess electrons must be added to an isolated spherical conductor 0.100 meters in diameter to produce an electric field of 1800 N/C just outside the surface?


1
Expert's answer
2022-04-04T09:07:21-0400

Answer

No of electrons added to sphere

"N=\frac{ER^2}{Ke}\\=\frac{1800*0.05*0.05}{9*10^9*1.6*10^{-19}}\\=31*10^6"